Consider a wire with resistivity of 2.20x10-7 Ω m, a diameter of 9.50 mm, and a length of 29 m. Determine the resistance of the wire in ohms (Ω). Round your answer to three significant figures.
Answer #1

Solution :

The resistance of the wire is given by R = \rho*l/a

where \rho is resistivity , l is the length of the wire and a is the area of cross section.

Considering the cylindrical wire : area of cross section of wire is  

a = \pi*(r^2) = 3.14*9.5*9.5*(10^(-6))/4

a = 7.085 * 10^(-5) m^2

Resistance : R = \rho*l/a

R = 2.2* 10^(-7)* 29 / 7.085* 10^(-5) = 9.005 * 10^(-2) ohm